# Quantifying community evolution in temporal networks

### v2

@@ IV DISCUSSION
The introduction of Union-NMI (UNMI) and Intersection-NMI (INMI) addresses a significant gap in network analysis, particularly in the comparison of network partitions with differing node sets. Our experiments on both synthetic and real-world temporal networks demonstrate that these metrics offer robust, flexible, and meaningful assessments of partition similarity where traditional metrics fall short. One of the key findings from our experiments is the effectiveness of UNMI and INMI in handling dynamic changes inherent in temporal networks. Traditional NMI assumes identical node sets between partitions, which is often not the case in real-world scenarios where nodes can join or leave the network over time. By explicitly considering the union or intersection of node sets, UNMI and INMI accommodate these changes without resorting to artificial adjustments that can bias results.
@@
One important aspect of our study is the interpretability of UNMI and INMI. By building upon the foundational concepts of mutual information and adjusting them to account for node-set differences, these metrics retain the intuitive appeal of NMI while extending its applicability. This makes them accessible tools for researchers who are already familiar with traditional network analysis techniques. Both UNMI and INMI are necessary and one should not be preferred over the other as they give different insights, particularly when data sets have high node churn. In some circumstances, the answers they give are broadly the same but when the node overlap between two different windows is small compared with the network size, the difference between the two measures can be large.
@@
A surprise in our study was that none of the data sets we looked at showed long-lasting community structures which persisted over time. The networks studied rarely showed evidence for strong similarity in community structure lasting for any long proportion of the data set. This is somewhat surprising, especially for a network like the email-EU-core network which is the same group of people interacting over time. While our sample of five studied networks is too small to draw very general conclusions, it brings into question studies which have considered the communities formed by observing the networks statically with all these time windows compressed into one static graph structure.
@@
In conclusion, our proposed metrics, UNMI and INMI, enhance the toolkit available for network partition comparison by directly addressing the challenges posed by differing node sets. They offer a balance of theoretical rigour, practical applicability, and computational efficiency. There are practical considerations and limitations to keep in mind (obviously it does require more computational power than static analysis), however, the benefits they provide beyond simply looking at a static network are considerable.

### v3

@@ 5 Discussion
The introduction of Union-AMI (UAMI) and Intersection-AMI (IAMI) addresses a significant gap in network analysis, particularly in the comparison of network partitions with differing node sets. Our experiments on both synthetic and real-world temporal networks demonstrate that these metrics offer robust, flexible, and meaningful assessments of partition similarity where traditional metrics fall short. One of the key findings from our experiments is the effectiveness of UAMI and IAMI in handling dynamic changes inherent in temporal networks. Traditional extension of MI assumes identical node sets between partitions, which is often not the case in real-world scenarios where nodes can join or leave the network over time. By explicitly considering the union or intersection of node sets, UAMI and IAMI accommodate these changes without resorting to artificial adjustments that can bias results.
@@
One important aspect of our study is the interpretability of UAMI and IAMI. By building upon the foundational concepts of mutual information and adjusting them to account for node-set differences, these metrics retain the intuitive appeal of mutual information while extending its applicability. This makes them accessible tools for researchers who are already familiar with traditional network analysis techniques. Both UAMI and IAMI are necessary and one should not be preferred over the other as they give different insights, particularly when data sets have high node churn. In some circumstances, the answers they give are broadly the same but when the node overlap between two different windows is small compared with the network size, the difference between the two measures can be large.
@@
A surprise in our study was that none of the data sets we looked at showed long-lasting community structures which persisted over time. The networks studied rarely showed evidence for strong similarity in community structure lasting for any long proportion of the data set. This is somewhat surprising, especially for a network like the email-EU-core network which is the same group of people interacting over time. While our sample of five studied networks is too small to draw very general conclusions, it brings into question studies which have considered the communities formed by observing the networks statically with all these time windows compressed into one static graph structure. Of course this conclusion depends on the community detection algorithms used and also on the size of the persistent communities. It is possible that smaller stable communities exist within these networks but this stability is not reflected by the partition of the whole network into communities.
@@
In conclusion, our proposed metrics, UAMI and IAMI, enhance the toolkit available for network partition comparison by directly addressing the challenges posed by differing node sets. They offer a balance of theoretical rigour, practical applicability, and computational efficiency. There are practical considerations and limitations to keep in mind (obviously it does require more computational power than static analysis), however, the benefits they provide beyond simply looking at a static network are considerable.
Author contributions statement
Peijie Zhong, Raúl Mondragón and Richard Clegg designed the methods involved in this article. Peijie Zhong performed the experiments and analysed the data. Cheick Ba, Raúl Mondragón and Richard Clegg advised on the analysis of the experiment data.
Data availability
The email-EU-core, Math Overflow, arXiv HEP-TH citation network and subReddit hyperlink network datasets are available at https://snap.stanford.edu/data/index.html and the NFT trading data is available at https://osf.io/wsnzr/?view_only=319a53cf1bf542bbbe538aba37916537.
Code availability
The code is available at https://github.com/Peijie-Zhong/community-similarity. The software Raphtory we used can be found at https://github.com/Pometry/Raphtory, the version number is v0.15.1.
Acknowledgement
Thanks to Pometry for giving help and support with their Raphtory software that was used for many results in this paper.
Additional information
The authors declare no competing interests.